Find y(t) solution of the initial value problem 2 y2 +bt2 y= y(1) = 1, t>0, ty
Answer #1

112) Hene y = 292 1 bt2 ; Y(1) =Z,, t>0 ty i dy dy = 2외자 642 dt ty 292 bt2 ty : - 2 + b Let Fly, t) LI 32 IT 12(월) 2(53)746 (Substitue all values wc get, V 11 t. do 11 vtt du 202 + b dt V t du 222 + b dt t. do = 2 v2 + b-v2 dt v2 tb dt V dv = at 22b.) ² + b = 2² 2² y² +bt2 = +42 - (1) we get Substitute t = 1, y = 1 in ean (1) 11) ²+ b(2)2 = (2)4 (2 1 + b = c2 Substitute c